A rectangular tank has a base 60cm by 20cm. A solid pyramid with a square base of sides 10cm each and height 27cm is placed inside the tack. The tank is than filled with water until it just covers the pyramid. If the pyramid is removed calculate the fall in the level of water in the tank.
Answer by Cromlix(4381)

Volume of pyramid = length * width of base * height *1/3
Volume of pyramid = 10 * 10 * 27 * 1/3 = 900 cm^3
Volume of tank required to just cover pyramid
= length * width of base * height (pyramid)
= 60 * 20 * 27 = 32400 cm^3
......................
the pyramid is now removed;
32400 - 900 = 31500 cm^3
.......................
New height of liquid:-
31500 cm^3 = 60 * 20 * height
height = 31500/(60 * 20)= 26.25 cm
A fall of 0.75 cm.
Hope this helps.
;-)